OpenStudy (anonymous):

What is the missing exponent? (12^-5)^2 l_____l ----------=12 12^-4

OpenStudy (anonymous):

@KJSaif

OpenStudy (anonymous):

help

Nnesha (nnesha):

you need to know 2 exponent rules \[\huge\rm (x^m)^n=x^{m \times n}\] and the 2nd one \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\]when we divide same bases we should subtract their exponents

Nnesha (nnesha):

so \[\huge\rm (12^{-5})^2 = ?\]

OpenStudy (anonymous):

61917364224

Nnesha (nnesha):

use the exponent rule i gave you don't use calculator .

Nnesha (nnesha):

so \[\huge\rm (12^{-5})^2 = ?\] \[\huge\rm (x^m)^n=x^{m \times n}\]multiply the exponents

OpenStudy (anonymous):

20^?

OpenStudy (anonymous):

i mean 12^?

Nnesha (nnesha):

what are the exponents \[\huge\rm (12^{-5})^2 = ?\] ?

OpenStudy (anonymous):



Nnesha (nnesha):

example \[\huge\rm (2^3)^4=2^{3 \times 4}=2^{12}\]

OpenStudy (anonymous):

so 12 ^2

Nnesha (nnesha):

let's solve the numerator first , look at the example i gave you \[\huge\rm (12^{-5})^2 = ?\]

Nnesha (nnesha):

multiply the 2 with the exponent of -12

OpenStudy (anonymous):

24

OpenStudy (anonymous):

144

Nnesha (nnesha):

\[\huge\rm (12^{-5})^2 = 12^{-5 \times 2}= 12^{??}\]